OpenStudy (anonymous):

Solve the equation by factoring. 4x^2 + 12x = 0

OpenStudy (anonymous):

SO I factored it to 4(x+3x) = 0 And so now should I divide the left by 4 so that it's \[\frac{ x+3x }{ 4 }=0\] or does tht make no sense to do? (Because it doesn't make sense to me)

OpenStudy (anonymous):

oops this was a dumb question. sorry

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Do u still have a question?

OpenStudy (anonymous):

4x^2+12x=0 4x(x+12)=0 x=0 or -12

OpenStudy (anonymous):

x = 0, -3.

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

x=0 or -12

OpenStudy (anonymous):

U factored it incorrectly. Look at mine

OpenStudy (anonymous):

Factor out 4x not just 4

OpenStudy (anonymous):

Well the program for my homework says im right